AS6301(1L) AS6301(1L) NAME as6301−crossassemblerformicrocomputers SYNOPSIS as6301[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theas6301commandassemblestheinputfileintoatextoutputfilerepresentingtheprogram memoryforamicrocomputer. Options -pcpuIgnoredbythisprogram. -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforas6301Frankenstein Assembler(fileas6301.doc) NOTES Thereisonlyoneinputfile. 1